business
Problem 2-19 Consider the finear program Max 34 + 40 s.t. 1A+ 28s 8 1A28s 12 2A+ 18s 16 A, 82 0 or leave the box blank the model, enter 0 for a. Write the problem in standard form, For those boxes n which you must enter sueractive or neostive nmbers use a meus sign, (Example:-300) f you dot need the vanable A. S e S St Max s.t. s A+ S A+ A, B, St, Sa, S b. Select the correct graph that shows the optimal solution for the problem 0) 14 14 12- 12-
Ch. 2 Homework Calculator b. Select the correct graph that shows the optimal solution for the problem. AB 14- AB 14- (i) (0) (3) (3) 12- 12- 10- (1) (1) 10- Optimal solution A-4.8, B-6.4 Value-40 Optimal solution A-20/3, B-8/3 Value-30 2/3 2- (2) (2) 10 10 12 12 () AB 14- AB 14- (iv) 0) (3) 12- 12- 10- (1) 10- (1) 8- Optimal solution Check My Work 2 more Check My Work uses remaining
(in) AB (iv) 4B 14- 14 (3) 12- 12- 10- 10- 8- Optimal solution A-48, B-64 Walue 40 Optimal solution A-20/3, B-8/3 Value-30 2/3 Graph (i) c. What are the values of the three slack variables at the optimal solution? If required, round your answers to one decimal place. If no entry is required, enter "0" or leave the box blank. Si S2 - Sa -
Answer #1

G LPP is iven 3 A+4B Max Siti LA+ 2B & IA + 2BC12 13 16 2 A t ntadlucing Slack ruablia a ) woite PP we Can in stanch ard fimnopresoucts as dauatrons &(3 nuguorlitaa Now egn can n-inte21cept Mtercept intarcept n has n-intancept- O 4 (3 12 11 2 12 2 34fisx co-ordinates of B At 28-12 X2 2ATB=G 3 13= Q B = A + 12 A 291 frs ca-crdinatrs of R -Ar+2 8 & 42 B-(2 A+10M1 2 A=2 3 AtSlack in constaain( Now gisen by At 2 BtS 8 2x8 tt-8 20 3 3 4 3 28 S.3 $2 ly Sim Slack m 7 At28 + S 12 16 20 2 +Sz12 12-12 SO